how many pounds of peanuts at php 80.00 per pound should be added to 200 pounds of cashew nuts at php 200.00 per pound of the mixture is to be worth php 120.00 per pound?
Answer by mananth(16946)

--------price ---------------- quantity
cashew 200 ---------------- 200 lb
peanuts 80 ---------------- x lb
Mixture 120 ---------------- 200 + x lb

200*200 +80x=120(200+x)
40000+80x=24000 +120x
80x+(-120x)=24000 -(-40000)
-40 x = -16000
/-40
x=400
400 lb of peanuts